# Learning Network Representations with Disentangled Graph Auto-Encoder

### v1

@@ 6.6 Convergence Behavior and Complexity Analysis
Figure 9 shows the evolution of testing AUC for DVGA and VGAE. Similar results regarding AP are depicted in Figure 13 in Appendix. As observed from the figure, VGAE tends to easily fall into local optima, while our model exhibits greater stability, converging to higher peaks. Regarding computational complexity, Table 8 presents the time taken for training in each epoch. DVGA is slower on three citation network datasets compared to VGAE, mainly because of the computation of the graph with a complexity of $O(N^{2})$. Nevertheless, considering significant enhancement in performance, we believe these costs are worthwhile, particularly with the enhancement in computational capabilities and stability.
Table 8: Average training time (s) per epoch.
@@
Refer to captionRefer to caption
(a) Cora
Refer to captionRefer to caption
(b) CiteSeer
Refer to captionRefer to caption
(c) PubMed
Figure 9: The convergence behavior of VGAE and our proposed DVGA on the test set for the link prediction task in terms of AUC. (a) Cora. (b) CiteSeer. (c) PubMed.
7 Conclusion
In this article, we investigate the problem of learning disentangled network representations using generative models and propose the Disentangled Graph Auto-Encoder (DGA) and Disentangled Variational Graph Auto-Encoder (DVGA) models. We design a dynamic disentangled graph encoder capable of aggregating features in a disentangled manner. Additionally, we introduce component-wise flow for the DVGA, encouraging DVGA to learn expressive representations that best describe graphs. We also propose a factor-wise decoder tailored for disentangled representations to improve its ability to predict graph structures. Independent regularization is introduced to eliminate statistical dependencies between different latent representations. Empirical results including quantitative and qualitative experiments on synthetic and real-world datasets demonstrate the effectiveness of our method in learning interpretable disentangled representations. As for future work, exploring new applications that leverage the robustness and interpretability brought by disentangled representations will be an interesting and promising research direction.
Acknowledgments
This work was funded by the National Nature Science Foundation of China under Grant No. 12320101001 and 12071428.
Appendix
(a) Number of channels $K$.
(b) Regularization $\lambda$
(c) Number of Layers $L$
Figure 10: Analysis of different hyper-parameters in terms of AP on Cora.
(a) Number of channels $K$.
(b) Regularization $\lambda$
(c) Number of Layers $L$
Figure 11: Analysis of different hyper-parameters in terms of AP on CiteSeer.
(a) Number of channels $K$.
(b) Regularization $\lambda$
(c) Number of Layers $L$
Figure 12: Analysis of different hyper-parameters in terms of AP on PubMed.
Refer to captionRefer to caption
(a) Cora
Refer to captionRefer to caption
(b) CiteSeer
Refer to captionRefer to caption
(c) PubMed
Figure 13: The convergence behavior of VGAE and our proposed DVGA on the test set for the link prediction task in terms of AP. (a) Cora. (b) CiteSeer. (c) PubMed.

### v2

@@ VI-F Convergence Behavior and Complexity Analysis
Figure 9 shows the evolution of testing AUC for DVGA and VGAE on Cora. Similar trends can be observed on Citeseer and PubMed. The results of testing AP exhibit a similarity to the AUC results, which we have excluded here for simplicity. As observed in the figure, VGAE tends to easily fall into local optima, while our model exhibits greater stability, converging to higher peaks. Regarding computational complexity, Table IX presents the time taken for training in each epoch. DVGA is slower on three citation network datasets compared to VGAE, mainly because of the computation of the graph with a complexity of $O(N^{2})$. Nevertheless, considering the significant enhancement in performance, we believe these costs are worthwhile, particularly with the enhancement in computational capability and stability.
Figure 9: The convergence behavior of our proposed DVGA and VGAE on Cora for the link prediction task in terms of AUC.
TABLE IX: Average training time (s) per epoch.
@@
VII Conclusion
In this article, we investigated the problem of learning disentangled network representations using the (variational) graph auto-encoder and proposed two models, namely Disentangled Graph Auto-Encoder (DGA) and Disentangled Variational Graph Auto-Encoder (DVGA). We developed a dynamic disentangled graph encoder that can effectively aggregate features in a disentangled manner. The utilization of component-wise flow in the DVGA further facilitates the learning of expressive representations. In order to better capture distinct information between different channels in latent representations, we implemented independent regularization to promote statistical independence. In addition, a factor-wise decoder tailored for disentangled representations was built to improve the models’ performance. Extensive experimental results on synthetic and real-world datasets demonstrated that our method outperforms recent baselines and effectively learns interpretable disentangled representations. As for future work, it would be worthwhile and intriguing to explore new applications that make use of the strength provided by disentangled representations.